OpenStudy (k8lyn911):

Calculus II - Logarithms Express the given quantity as a single logarithm. (1/5) ln (x+2)^(5) + (1/2) [ln(x)-ln(x^(2)+3x+2)^(2)]

OpenStudy (jhannybean):

\[\frac{1}{5}\ln(x+2)^5 +\frac{1}{2}[\ln(x)-\ln(x^2+3x+2)^2]\]natural logarithms work the same way as log functions, same rules, so \[\ln(x^2) \iff 2\ln(x)\]\[\ln(a) +\ln(b) = \ln(a \cdot b)\]

OpenStudy (jhannybean):

And always remember what your main function is, that would be the addition.

OpenStudy (jhannybean):

\[\frac{1}{5}\ln(x+2)^5 \implies \left[\ln(x+2)^5\right]^{1/5} \implies ~?\]

OpenStudy (k8lyn911):

5th root ( ln ((x+2)^(5)) ?

OpenStudy (jhannybean):

What is \(\large (a^5)^{1/5} = ~?\)

OpenStudy (k8lyn911):

a

OpenStudy (jhannybean):

Same rule applies to this: \(\color{blue}{\text{Originally Posted by}}\) @Jhannybean \[\frac{1}{5}\ln(x+2)^5 \implies \left[\ln(x+2)^5\right]^{1/5} \implies ~?\] \(\color{blue}{\text{End of Quote}}\) Think of \(\ln(x+2)\) as \(a\).

OpenStudy (k8lyn911):

Is it just ln(x+2) then?

OpenStudy (jhannybean):

Yep \(\checkmark\)

OpenStudy (jhannybean):

Now for \(\frac{1}{2}[\ln(x) -\ln(x^2+3x+2)^2]\) We could distribute the \(\frac{1}{2} t both terms inside the brackets. What would we get?

OpenStudy (jhannybean):

\(\frac{1}{2}\)*

OpenStudy (k8lyn911):

(1/2) ln (x) - (1/2) ln ((x^(2) + 3x + 2)^(2)) = ln ( sqrt (x)) - ln ((x^(2) + 3x + 2)^(2))^(1/2) = ln ( sqrt (x)) - ln (x^(2) + 3x + 2)

OpenStudy (jhannybean):

Alright, so you have \[\ln(\sqrt{x})-\ln(x^2+3x+2) \iff \ln(x^{1/2}) -\ln(x^2+3x+2)\]Which is correct \(\checkmark\)

OpenStudy (k8lyn911):

So using the properties of logs, ln (x+2) + ln (sqrt(x)) - ln (x^(2) + 3x + 2) would be ln ( ((x+2)(sqrt(x))/((x)^(2)+3x+2)) ) ?

OpenStudy (k8lyn911):

\[\ln(\frac{(x+2)\sqrt{x}}{ x^2 + 3x + 2 })\]

OpenStudy (jhannybean):

That's correct :)

OpenStudy (jhannybean):

\[\log(a) + \log(b) - \log(c) \iff \log\left(\frac{ab}{c}\right)\]

OpenStudy (jhannybean):

But now you need to simplify \(x^2+3x+2\)

OpenStudy (k8lyn911):

\[\ln(\frac{ \sqrt{x}(x+2) }{ (x+1)(x+2) }) = \ln(\frac{ \sqrt{x} }{ x+1 })\]

OpenStudy (jhannybean):

Ohh haha my bad. Sorry.

OpenStudy (k8lyn911):

Oh. Our teacher prefers using ln.

OpenStudy (jhannybean):

yeah, I just misread my own post, lol.

OpenStudy (k8lyn911):

Thanks for explaining everything. :)

OpenStudy (jhannybean):

But that's correct. \[\ln\left(\frac{x^{1/2}}{x+1}\right) \iff \ln(x^{1/2}) -\ln(x+1) \iff \frac{1}{2}\ln(x) - \ln(x+1)\]

OpenStudy (jhannybean):

No problem :)